# Alias-Free Latent Diffusion Models: Improving Fractional Shift Equivariance of Diffusion Latent Space

### v1

@@ Abstract
Figure 1: Visualization of AE and LDM latent shift experiments. Top: we shift the input of an 8×8$\times$8 × upsampling VAE decoder by 1/8 pixel in each step. The intermediate output images from Stable Diffusion (SD) VAE appear increasingly blurry. Bottom: we shift the noisy latent of LDM by 1/8 pixel in each step. Without cross-frame attention, SD struggles to generate consistent results; even with it, we can observe a “bouncing effect”, where textures initially stick but suddenly shift in subsequent steps. In contrast, our alias-free SD (AF-SD) produces shift-equivariant results in both experiments, maintaining consistency and clarity across steps.
@@ 1 Introduction
Figure 2: Frequency map visualization of upsampled VAE latent. The upsampled latent serves as an approximation of a continuous signal. Left: Illustration of latent upsampling algorithm. For an encoder that downsamples the image by 8×8$\times$8 ×, we can leverage its shift-equivariance to achieve 2×2$\times$2 × upsampling of the latent. By shifting the input image by 4 pixels, we create a 0.5-pixel shift in the latent. Combining these fractional shifted latents yields a higher-resolution latent. Right: Visualization of 8×8$\times$8 × upsampled latent in both spatial and frequency domains. (a) AF-VAE with random weights, (b) AF-VAE without equivariance loss, (c) AF-VAE with equivariance loss. Ideally, a correctly band-limited latent would only use $1/8$ of the frequencies in the upsampled frequency domain. During training, the aliasing effects in a randomly initialized VAE with alias-free modules become more prominent; however, our proposed equivariance loss helps suppress the aliasing during training.
@@ 2.2 Stability of Latent Diffusion Models
Compared to GANs, LDMs exhibit notable instability, especially concerning smooth latent space and shift-equivariance. A smooth latent space, a concept well-developed in GAN literature, enables continuous changes in generated images when interpolating latents [16, 17]. This smoothness enhances image editing applications for diffusion models, such as inversion [43, 31] and interpolation [54, 40, 13]. The other property, shift-equivariance, ensures consistency in generated outputs as input objects shift within a scene. This property is crucial for achieving temporal coherence in video editing applications [21, 50, 51, 5, 10, 4]. Common practices that mitigate the instability of diffusion models include image overfitting [19, 41, 54], cross-frame attention [21], and regularization [11, 52]. However, the shift-equivariance of LDM is still an underexplored topic.
@@ 3.1 Preliminaries
StyleGAN3 StyleGANs [16, 17] generate high-resolution images by gradually upsampling a constant low-resolution feature with a CNN-based synthesis network. Despite high generation quality, a “texture sticking” effect occurs in the original StyleGAN, where the details in the output image stick to the same position when shifting the input feature. StyleGAN3 [18] argues that the aliasing in the synthesis network causes this effect. The issue is analyzed by a continuous signal interpretation: each network feature is equivalent to a continuous signal. A discrete feature $Z$ can be converted into a continuous signal $z$ via interpolation filter and be converted back via Dirac comb. Similarly, each network layer $F$ has its counterpart $f$ in the continuous domain. If a network layer is shift-equivariant, then its effect in the discrete and continuous domain should be equivalent, i.e., the conversion between $F(Z)$ and $f(z)$ should still be invertible.
@@
4. Filtered Nonlinearity: Wrap nonlinearities between a 2×2$\times$2 × upsampling and 2×2$\times$2 × downsampling to suppress the high frequencies introduced by this operation.
@@ 3.3 Continuous Latent Representation via Equivariance Loss
$L_{\text{enc}}=||\left[\mathcal{E}(T_{\Delta}(x))-T_{\Delta/k}(\mathcal{E}(x))\right]\cdot M_{\Delta/k}||_{2}^{2},$ (4)
$L_{\text{dec}}=||\left[\mathcal{D}(T_{\Delta/k}(z))-T_{\Delta}(\mathcal{D}(z))\right]\cdot M_{\Delta}||_{2}^{2},$ (5)
@@
$L_{\text{unet}}=||\left[\epsilon^{\prime}_{\theta}(T_{\Delta}^{\text{cir}}(z_{t}),t)-T_{\Delta}(\epsilon_{\theta}(z_{t},t))\right]\cdot M_{\Delta}||_{2}^{2},$ (6)
@@ 4.1 Ablation Study
The quantitative results are shown in Table 3, and partial visualization results are shown in Fig. 4. Although LDM with SD VAE has good SPSNR, we can observe severe quality degradation in fractional shift results (Fig. 4(a)). Using AF-VAE can improve the quality of shifted frames, but “texture sticking” still occurs (Fig. 4(b)). This issue can be addressed by adding equivariance loss (Fig. 4(c)). By applying CFA in equivariance loss, the final model, AF-LDM, has the best overall shift-equivariance.
@@ 4.3 Image-to-image Translation
Super-resolution with Latent I2SB. Image-to-Image Schrödinger Bridge (I2SB) [25] is a conditional diffusion model that directly maps one image distribution to another without going through Gaussian noise. Originally, I2SB operates in image space, with model weights initialized from an unconditional diffusion model. In our experiments, we implement two variants of I2SB for 4×4$\times$4 × super-resolution in latent space: one based on LDM and the other on AF-LDM. As shown in Fig. 9, the standard latent I2SB exhibits severe flickering and quality degradation when input shifts. In contrast, the alias-free variant demonstrates strong shift-equivariance, providing consistent outputs.
Figure 9: Qualitative comparison of shift-equivairance in latent I2SB 4×4$\times$4 × super-resolution results.
@@ A.1 Alias-free Modules
Downsampling. Strided downsampling convolutions are replaced with a standard convolution followed by a low-pass filter and nearest downsampling. The low-pass filter is implemented as an ”ideal low-pass filter,” removing high frequencies in the Fourier domain, as described in [30].
Upsampling. Upsampling is performed by zero-padding between existing pixels followed by convolution with a sinc interpolation kernel. Similar to the “ideal low-pass filter” in downsampling, this convolution is executed in the Fourier domain via multiplication.
Nonlinearity. All nonlinearities, except those in the last layer, are wrapped between a 2×2$\times$2 × ideal upsampling and a 2×2$\times$2 × ideal downsampling.
@@ A.2 VAE
Resolution $256\times 256$ $256\times 256$ $512\times 512$
@@
Learning Rate $5\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-6}$
@@
where $x\in\mathbb{R}^{H\times W\times 3}$ is the input image, $z=sg(\mathcal{E}(x)),z\in\mathbb{R}^{H/k\times W/k\times 4}$ is the latent downsampled by $k\times$, $sg$ is stop gradient operator, and $M_{\Delta}$ denotes the valid mask for cropped shift $T_{\Delta}$. Offsets $\Delta=(\Delta_{x},\Delta_{y})\in\mathbb{N}^{2}$. $\Delta_{x},\Delta_{y}$ are uniformly sampled from $[-\frac{3}{8}H,\frac{3}{8}H]$ and $[-\frac{3}{8}W,\frac{3}{8}W]$, respectively.
@@ A.3 Unconditional LDM
$L^{\text{LDM}}=\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,1),t},[||\epsilon-\epsilon_{\theta}(z_{t},t)||_{2}^{2}]+\lambda L_{\text{eq}}^{\text{LDM}},$ (15)
@@
$L_{\text{eq}}^{\text{LDM}}=\mathbb{E}_{\mathcal{E}(x),t}||\left[\epsilon^{\prime}_{\theta}(T_{\Delta}^{\text{cir}}(z_{t}),t)-T_{\Delta}(\epsilon_{\theta}(z_{t},t))\right]\cdot M_{\Delta}||_{2}^{2},$ (16)
@@ A.4 Text-conditional LDM
$L^{\text{SD}}=\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathbf{N}(0,1),t,c},[||\epsilon-\epsilon_{\theta}(z_{t},t,c)||_{2}^{2}]+\lambda L_{\text{eq}}^{\text{LDM}},$ (17)
@@ B Details of Video Editing
3:$z^{1}_{T}\leftarrow\texttt{DDIMInversion}(z^{1}_{0},\epsilon_{\theta})$triangleright Cache K, V in $\epsilon^{\prime}_{\theta}$
4:for $i$ in $[2,...,N]$ do
@@
6:$\hat{z}^{1}_{0}\leftarrow\texttt{DDIMSampling}(z^{1}_{T},\epsilon_{\theta})$triangleright Cache K, V in $\epsilon^{\prime}_{\theta}$
7:$\hat{x}^{1}\leftarrow\mathcal{D}(\hat{z}^{1}_{0})$
8:for $i$ in $[2,...,N]$ do
@@
1:$f_{fwd},f_{bwd}\leftarrow\texttt{FlowEstimation}(x^{1},x^{2})$
2:$z^{1}_{0}\leftarrow\mathcal{E}(x^{1})$
3:$z^{2}_{0}\leftarrow\mathcal{E}(x^{2})$
4:$z^{1}_{T}\leftarrow\texttt{DDIMInversion}(z^{1}_{0},\epsilon_{\theta})$
5:$z^{2}_{T}\leftarrow\texttt{DDIMInversion}(z^{2}_{0},\epsilon_{\theta})$
6:$\texttt{DDIMSampling}(z^{1}_{0},\epsilon_{\theta})$triangleright Cache K, V in $\epsilon^{\prime 1}_{\theta}$
7:$\texttt{DDIMSampling}(z^{2}_{0},\epsilon_{\theta})$triangleright Cache K, V in $\epsilon^{\prime 2}_{\theta}$
8:for $i$ in $[1,...,N]$ do
@@ B.2 Image Splatting and Interpolation
Leveraging AF-SD’s high warping-equivariance, we also implement a simple image interpolation method using latent splatting (Alg. 2). Given two input images, we invert their latents, splat them using forward and backward flows with interpolation ratio $\alpha$, and interpolate the latents and attention features to generate intermediate frames [54, 13]. Compared to standard SD, AF-SD produces smoother results. Although the method sometimes produces flickering artifacts in the occlusion region caused by latent splatting, this issue can be mitigated by integrating depth information and softmax splatting [32]. To keep the algorithm as simple as possible, we do not apply this more advanced technique. The results are shown in Fig. 12.

### v2

@@ Abstract
Figure 1: Visualization of AE and LDM latent shift experiments. Top: we shift the input of an $8\times$ upsampling VAE decoder by 1/8 pixel in each step. The intermediate output images from Stable Diffusion (SD) VAE appear increasingly blurry. Bottom: we shift the noisy latent of LDM by 1/8 pixel in each step. Without cross-frame attention, SD struggles to generate consistent results; even with it, we can observe a “bouncing effect”, where textures initially stick but suddenly shift in subsequent steps. In contrast, our alias-free SD (AF-SD) produces shift-equivariant results in both experiments, maintaining consistency and clarity across steps.
@@ 1 Introduction
Figure 2: Frequency map visualization of upsampled VAE latent. The upsampled latent serves as an approximation of a continuous signal. Left: Illustration of latent upsampling algorithm. For an encoder that downsamples the image by $8\times$, we can leverage its shift-equivariance to achieve $2\times$ upsampling of the latent. By shifting the input image by 4 pixels, we create a 0.5-pixel shift in the latent. Combining these fractional shifted latents yields a higher-resolution latent. Right: Visualization of $8\times$ upsampled latent in both spatial and frequency domains. (a) AF-VAE with random weights, (b) AF-VAE without equivariance loss, (c) AF-VAE with equivariance loss. Ideally, a correctly band-limited latent would only use $1/8$ of the frequencies in the upsampled frequency domain. During training, the aliasing effects in a randomly initialized VAE with alias-free modules become more prominent; however, our proposed equivariance loss helps suppress the aliasing during training.
@@ 2.2 Stability of Latent Diffusion Models
Compared to GANs, LDMs exhibit notable instability, especially concerning smooth latent space and shift-equivariance. A smooth latent space, a concept well-developed in GAN literature, enables continuous changes in generated images when interpolating latents [17, 18]. This smoothness enhances image editing applications for diffusion models, such as inversion [44, 32] and interpolation [56, 41, 14]. The other property, shift-equivariance, ensures consistency in generated outputs as input objects shift within a scene. This property is crucial for achieving temporal coherence in video editing applications [22, 52, 53, 5, 11, 4]. Common practices that mitigate the instability of diffusion models include image overfitting [20, 42, 56], cross-frame attention [22], and regularization [12, 54]. Warped Diffusion [6] specifically identifies the issue of shifting inconsistency in LDMs and proposes equivariance self-guidance as a remedy. Despite these efforts, the shift-equivariance of LDMs is still an underexplored topic.
@@ 3.1 Preliminaries
StyleGAN3 StyleGANs [17, 18] generate high-resolution images by gradually upsampling a constant low-resolution feature with a CNN-based synthesis network. Despite high generation quality, a “texture sticking" effect occurs in the original StyleGAN, where the details in the output image stick to the same position when shifting the input feature. StyleGAN3 [19] argues that the aliasing in the synthesis network causes this effect. The issue is analyzed by a continuous signal interpretation: each network feature is equivalent to a continuous signal. A discrete feature $Z$ can be converted into a continuous signal $z$ via interpolation filter and be converted back via Dirac comb. Similarly, each network layer $F$ has its counterpart $f$ in the continuous domain. If a network layer is shift-equivariant, then its effect in the discrete and continuous domain should be equivalent, i.e., the conversion between $F(Z)$ and $f(z)$ should still be invertible.
@@
4. Filtered Nonlinearity: Wrap nonlinearities between a $2\times$ upsampling and $2\times$ downsampling to suppress the high frequencies introduced by this operation.
@@ 3.3 Continuous Latent Representation via Equivariance Loss
[…]
@@
[…]
@@ 4.1 Ablation Study
The quantitative results are shown in Table 3, and partial visualization results are shown in Fig. 4. Although LDM with SD VAE has good SPSNR, we can observe severe quality degradation in fractional shift results (Fig. 4(a)). Using AF-VAE can improve the quality of shifted frames, but “texture sticking" still occurs (Fig. 4(b)). This issue can be addressed by adding equivariance loss (Fig. 4(c)). By applying CFA in equivariance loss, the final model, AF-LDM, has the best overall shift-equivariance.
@@ 4.3 Image-to-image Translation
Super-resolution with Latent I2SB. Image-to-Image Schrödinger Bridge (I2SB) [26] is a conditional diffusion model that directly maps one image distribution to another without going through Gaussian noise. Originally, I2SB operates in image space, with model weights initialized from an unconditional diffusion model. In our experiments, we implement two variants of I2SB for $4\times$ super-resolution in latent space: one based on LDM and the other on AF-LDM. As shown in Fig. 9, the standard latent I2SB exhibits severe flickering and quality degradation when input shifts. In contrast, the alias-free variant demonstrates strong shift-equivariance, providing consistent outputs.
Figure 9: Qualitative comparison of shift-equivairance in latent I2SB $4\times$ super-resolution results.
@@ A.1 Alias-free Modules
Downsampling. Strided downsampling convolutions are replaced with a standard convolution followed by a low-pass filter and nearest downsampling. The low-pass filter is implemented as an "ideal low-pass filter," removing high frequencies in the Fourier domain, as described in [31].
Upsampling. Upsampling is performed by zero-padding between existing pixels followed by convolution with a sinc interpolation kernel. Similar to the “ideal low-pass filter" in downsampling, this convolution is executed in the Fourier domain via multiplication.
Nonlinearity. All nonlinearities, except those in the last layer, are wrapped between a $2\times$ ideal upsampling and a $2\times$ ideal downsampling.
@@ A.2 VAE
Resolution $\times 256 256×256$$\times 256$ $512\times 512$
@@
Learning Rate $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-6}$
@@
where $x\in\mathbb{R}^{H\times W\times 3}$ is the input image, $z=sg(\mathcal{E}(x)),z\in\mathbb{R}^{H/k\times W/k\times 4}$ is the latent downsampled by $k\times$, $sg$ is stop gradient operator, and $M_{\Delta}$ denotes the valid mask for cropped shift $T_{\Delta}$. Offsets $\Delta=(\Delta_{x},\Delta_{y})\in\mathbb{N}^{2}$. $\Delta_{x},\Delta_{y}$ are uniformly sampled from $\frac{3}{8}H,\frac{3}{8}H]$ and $\frac{3}{8}W,\frac{3}{8}W]$, respectively.
@@ A.3 Unconditional LDM
[…]
@@
[…]
@@ A.4 Text-conditional LDM
[…]
@@ B Details of Video Editing
[…]
4:for $i$ in [2,,N][2,...,N] do
@@
[…]
8:for $i$ in [2,,N][2,...,N] do
@@
[…]
8:for $i$ in [1,,N][1,...,N] do
@@ B.2 Image Splatting and Interpolation
Leveraging AF-SD’s high warping-equivariance, we also implement a simple image interpolation method using latent splatting (Alg. 2). Given two input images, we invert their latents, splat them using forward and backward flows with interpolation ratio $\alpha$, and interpolate the latents and attention features to generate intermediate frames [56, 14]. We use GMFlow [51] as the flow estimator. Compared to standard SD, AF-SD produces smoother results. Although the method sometimes produces flickering artifacts in the occlusion region caused by latent splatting, this issue can be mitigated by integrating depth information and softmax splatting [33]. To keep the algorithm as simple as possible, we do not apply this more advanced technique. The results are shown in Fig. 12.